OpenStudy (anonymous):

Consider the following sets of numbers. A={all +ve even numbers less than 25} B={all prime numbers less than 25} Suppose a number is chosen at random from 1 to 25. (a) Find P(A), P(B) and P(A∩B). (b) Hence find (AUB). @marigirl

OpenStudy (anonymous):

@marigirl is it include 25?

OpenStudy (marigirl):

no i will not include 25 .. anyway it wont matter P(A)= total of 12 numbers. therefore if a random number is chosen from 1-25, what will be the probability it will be a even number P(A)=12/25

OpenStudy (marigirl):

do u have ans to check

OpenStudy (anonymous):

P(A)=12/25

OpenStudy (marigirl):

awesome. i checked out and found out there are 9 prime numbers for our given restiction. so what will P(B) be?

OpenStudy (anonymous):

1 isn't prime number, is it?

OpenStudy (marigirl):

no, 1 is not prime (plus i goggled it :P)

OpenStudy (anonymous):

2, 3, 5, 7, 11, 13, 17, 19, 23 is 2 a prime number?

OpenStudy (marigirl):

2, 3, 5, 7, 11, 13, 17, 19, 23 yes ur on track

OpenStudy (anonymous):

9/25

OpenStudy (marigirl):

yes

OpenStudy (anonymous):

then P(A∩B)=1/25

OpenStudy (anonymous):

2, 3, 4, 5, 6, 7, 8, 10, 11, 12, 13, 14, 16, 17, 18, 19, 20, 22, 23 ,24

OpenStudy (marigirl):

sorry for the late reply

OpenStudy (marigirl):

I agree with that: then P(A∩B)=1/25

OpenStudy (marigirl):

last one Hence find (AUB).

OpenStudy (anonymous):

20/25=4/5

OpenStudy (marigirl):

P(A U B) = P(A) + P(B)

OpenStudy (anonymous):

and minus 1

OpenStudy (marigirl):

but it means everything that is in either of the sets

OpenStudy (anonymous):

but '2' is repeated

OpenStudy (marigirl):

no ur right, sorry

OpenStudy (marigirl):

thats why they wrote "Hence" find A U B

OpenStudy (marigirl):

anyways i gotta go, well done, you did most of the work!

OpenStudy (anonymous):

thank you haha